1. in interval notation, what is the range of F(x)=2x^3-4x-7 (i'm thinking that this range cannot be determined, am i correct.)
2. which of the following is a line perpendicular to y=_x+4 y=_x-_, y=2x-4,y=2x-_, or y=-2x
3. what is the slope of the line that passes through(-4,-2) and the origin
M=_, B=_, M=-_ or insufficient information given
4. two lines are parallel if m1=(-m2)-1, m1=-m2, m1=m2, or none of the above (i'm thinking none of the above is this correct)
5.f(x+h)-f(x) is this the numerator of the different quotient

Answer by stanbon(75887)

1. in interval notation, what is the range of F(x)=2x^3-4x-7 (i'm thinking that this range cannot be determined, am i correct.)
Graph it and you will see the range is "all Real Numbers".
Interval notation: (-inf,+inf)
----------------------------------------
2. which of the following is a line perpendicular to y=_x+4 y=_x-_, y=2x-4,y=2x-_, or y=-2x
I can't make out your equations.
-----------------------------------------
3. what is the slope of the line that passes through(-4,-2) and the origin
M=_, B=_, M=-_ or insufficient information given
slope = -2/4 = -1/2
--------------------------------------
4. two lines are parallel if m1=(-m2)-1, m1=-m2, m1=m2, or none of the above (i'm thinking none of the above is this correct)
Ans: m1=m2
-------------------------------
5.f(x+h)-f(x) is this the numerator of the different quotient
Ans: yes
========================
Cheers,
Stan H.